Find all real or imaginary solutions.
3v^2 + 4v - 1 = 0

Textbook Elementary and Intermediate Algebra. Mark Dugopolski
ISBN: 13978-0-07-326862-0

Answer by mukhopadhyay(490)

3v^2 + 4v - 1 = 0
=>v=[(-4)+sqrt(16+12)]/6 or v=[(-4)-sqrt(16+12)]/6 (using quadratic formula)
=>v=[(-4)+2sqrt(7)]/6 or v=[(-4)-2sqrt(7)]/6
=>v=[(-2)+sqrt(7)]/3 or v=[(-2)-sqrt(7)]/3